.15x-0.9(x-1)=0.01(3x-3) I can't figure out how to solve linear equation using decimals.
Found 2 solutions by CubeyThePenguin, Alan3354:
Answer by CubeyThePenguin(3113)

0.15x - 0.9(x - 1) = 0.01(3x - 3)
0.15x - 0.9x + 0.9 = 0.03x - 0.03
-0.75x + 0.9 = 0.03x - 0.03
0.87 = 0.78x

x = (0.87)/(0.78) = 87/78

Answer by Alan3354(69443)

.15x-0.9(x-1)=0.01(3x-3)
----
Multiply by 100 to eliminate the decimals
15x-90(x-1) = (3x-3)